# LIO-GVM: an Accurate, Tightly-Coupled Lidar-Inertial Odometry with Gaussian Voxel Map

### v1

@@ Abstract
This letter presents an accurate and robust Lidar Inertial Odometry framework. We fuse LiDAR scans with IMU data using a tightly-coupled iterative error state Kalman filter for robust and fast localization. To achieve robust correspondence matching, we represent the points as a set of Gaussian distributions and evaluate the divergence in variance for outlier rejection. Based on the fitted distributions, a new residual metric is proposed for the filter-based Lidar inertial odometry, which demonstrates an improvement from merely quantifying distance to incorporating variance disparity, further enriching the comprehensiveness and accuracy of the residual metric. Due to the strategic design of the residual metric, we propose a simple yet effective voxel-solely mapping scheme, which only necessities the maintenance of one centroid and one covariance matrix for each voxel. Experiments on different datasets demonstrate the robustness and accuracy of our framework for various data inputs and environments. To the benefit of the robotics society, we open source the code at https://github.com/Ji1Xingyu/lio_gvm.
@@
LiDAR inertial odometry, SLAM, voxel map, distribution
@@ I Introduction
The vital requirement for an autonomous robot is to estimate its ego-motion and build a map of the environment without prior knowledge, which is widely known as Simultaneous Localization and Mapping (SLAM). In particular, the light detection and ranging (LiDAR) sensor is commonly employed for performing SLAM because of its ability to provide light-invariant and accurate range measurements. However, the LiDAR-only SLAM system may fail in degenerate scenarios (few constraints for optimization) or high-speed applications (motion distortion in the cloud). For complementary, researchers combine IMU with LiDAR [1, 2, 3, 4, 5, 6, 7, 8] to form a LiDAR Inertial Odometry (LIO), as IMU can provide accurate short-term motion information and is insensitive to surroundings, thus can provide an initial pose guess and compensate for the distortion of LiDAR. Filter-based LIO [9, 10, 6, 8, 7] has been investigated in the recent few years due to its temporal and computational efficiency. The performance of a filter-based LIO system is dominated by two parts: the front end for state estimation and the mapping module for map maintenance.
The front end of a LIO system usually consists of four steps: state prediction using IMU, correspondence matching, residual formulation, and state estimation. For correspondence matching, the common assumption of these filter-based LIO systems is that the point in the current scan can be approximately projected to the correct feature in the target map. This projection is dominated by the summation of historical state estimations and the state prediction using IMU. However, the accumulation of localization drifts in long-term odometry is inevitable; Simultaneously, the presence of spike noise (a severe issue in IMU, especially for the low-cost ones) leads to inaccuracies in the IMU pose guess. Thus, the assumption would be broken as shown in Fig. 1. To cope with these issues, we represent the points in both the source and target cloud as a set of Gaussian probability distributions [11, 12, 13] and evaluate the matched correspondences by the disparity in variance. Moreover, based on the distribution assumption, we propose a new residual metric for filter-based LIO, which allows it to assess not just the distance but also the divergence in variance. The new formula quantifies both the distance and variance disparity, further enriching the comprehensiveness and accuracy of the residual metric.
(a) the nearest search.
(b) the near and similar search.
Figure 1: Comparison of different search schemes. The blue and red points denote the point in current scan and target map respectively. The nearest search may encounter the mismatch problem, while the near and similar search can reject the pairs with large disparity in variance.
For the maintenance of the built map, k-d tree (and its variant) [14] is adopted for many systems [10, 8, 9] in the mapping module, as it can provide strict K nearest neighbor (k-NN) search for the correspondence matching. Some systems opt to embed a supplemental data structure (k-d tree, Hilbert Curve, etc.) within voxels to form a hierarchical structure [6, 7], which leverages the rapid operation capabilities of voxels while also facilitating points management. Instead of storing all the raw points, we propose an incremental voxel map without any additional structures, which only maintains a centroid and covariance matrix for each voxel. The idea of this simple map scheme is inspired by the following aspects: 1) Due to the strategic design of the residual metric, the map module only necessitates the maintenance of the centroid and covariance matrix for each distribution instead of the raw points. Similar to VGICP [13], we aggregate all the distributions within a voxel (multi-point distributions to a single-voxel distribution) to estimate the distribution for the voxel robustly. Thus, we can only maintain one centroid and one covariance matrix for each voxel. 2) The strict k-NN search is unnecessary for the residual formulation since we can access the voxel in the target map where the source point occupies in. Moreover, we can evaluate the adjacent voxels by the variance divergence. 3). The time complexity for the delete, insert, and query operation of voxels (implemented by a hash table) is $\mathcal{O}(1)$, which is considerably faster than the k-d tree.
@@
In conclusion, we implement the proposed residual metric and mapping scheme to a filter-based LIO system denoted as LIO-GVM. The contributions can be summarized as follows:
  • We propose a new residual metric for filter-based LIO, which can reject mismatched correspondences as well as improve accuracy. The metric demonstrates an improvement from merely quantifying distance to incorporating variance disparity, further enriching the comprehensiveness and accuracy of the residual formulation.
@@
  • We propose a simple yet efficient map scheme in which the voxel map only necessitates the maintenance of the centroid and covariance matrix for each voxel. Due to the strategic design of the residual metric, the map module can achieve fast operation while maintaining the accuracy of state estimation.
@@
  • We implement the two key techniques into a filter-based LIO system and evaluate its efficiency through various experiments. To the benefit of the robotics community, our implementation code is open-source at https://github.com/Ji1Xingyu/lio_gvm.
II RELATED WORKS
The SLAM problem was formulated as a state estimation problem in the early state, e.g., ESKF [15] and RBPF [16]. Many previous studies have thoroughly discussed the advantages and disadvantages of EKF [17, 18]. For example, EKF is real-time and fast, but its Markov property makes it hard to tackle long-time associations like large-scale loop closing. From the optimization perspective, EKF can also be regarded as a single-time, non-iterating factor graph with three factors (the prior, the motion, and the observation) that immediately marginalizes the previous state [19]. In LO or LIO algorithms, EKF is a prevalent backend estimator choice since the lidar point clouds are usually accurate enough to build a local map [20]. The performance of a LIO system is dominated by two factors: point cloud registration and map scheme.
For point cloud registration, LOAM [20] first proposes to extract the point and line features from the source and target cloud for fast registration and achieve accurate results. MULLS [21] extends the feature types and solves the optimization via a multi-metric linear least square method. Some systems only extract features in the target cloud and direct register the raw source cloud into the target map [10]. NDT [11] regards points in the target cloud as Gaussian distributions. GICP [12] fits the points in both the source and target cloud to Gaussian distributions. NDT tends to be more computationally efficient and GICP tends to be more accurate. VGICP [13] voxelized the distributions in the target map and utilizes the voxel-based association approach, which greatly increases the efficiency while retaining comparable accuracy with GICP. LiTAMIN2 [22] introduces symmetric KL-Divergence for the residual metric, which can achieve accurate pose estimation even with a small number of registered points.
For map schemes, a common method is to maintain the target map with k-d tree [20, 23, 24] for the nearest neighbor searching in the feature association phase. The conventional k-d tree needs to rebuild the entire tree for the update, which is inefficient and time-consuming. FAST-LIO2 uses the ikd-Tree [14], a dynamic k-d tree that can incrementally update the map.Faster-LIO [6] additionally introduces a pseudo-Hilbert curve inside the voxel for the acceleration of correspondence matching while maintaining accuracy.
@@ III PRELIMINARY
${}^{L_{k}}\mathbf{p}_{j}$ The $j_{th}$ point with coordinate reference to $L$ at time $t_{k}$
${}^{G}\mathbf{v}_{i}$ The coordinate of the $i_{th}$ voxel reference to $G$
@@ III-A Notation
The important notations used in this letter are listed in Tab. I. Besides, we assume a static calibrated extrinsic matrix ${}^{L_{k}}\mathbf{T}_{I_{k}}=({}^{L_{k}}\mathbf{R}_{I_{k}},{}^{L_{k}}\mathbf{p}_{I_{k}})$ from IMU frame to LiDAR frame.
@@ III-B System Description
where ${}^{G}\mathbf{v}^{T}_{I}$ and ${}^{G}\mathbf{g}$ are the velocity of IMU and gravity vector reference to the global frame, $\mathbf{b}^{T}_{\boldsymbol{\omega}}$ and $\mathbf{b}^{T}_{\mathbf{a}}$ are the bias vectors of gyroscope and accelerometer of the IMU.
III-C Iterative Error State Kalman Filter
In the system state $\mathbf{x}$ defined above, the rotation matrix ${}^{G}\mathbf{R}_{I}$ are defined on $SO(3)$ in an over-parametrized way to avoid singularity, which makes it non-straightforward to estimate the system state on the manifold directly. An alternative way to estimate the state is to divide the true state $\mathbf{x}$ into two states defined on different spaces:
@@ III-C Iterative Error State Kalman Filter
$\delta\mathbf{x}_{k}$ $=\begin{bmatrix}{}^{G}\delta\mathbf{r}_{I_{k}}^{T}&{}^{G}\delta\mathbf{p}_{I_{k}}^{T}&{}^{G}\delta\mathbf{v}^{T}_{I_{k}}&\delta\mathbf{b}^{T}_{\boldsymbol{\omega}_{k}}&\delta\mathbf{b}^{T}_{\mathbf{a}_{k}}&{}^{G}\delta\mathbf{g}_{k}^{T}\end{bmatrix}^{T},$
@@ IV METHODOLOGY
The workflow of LIO-GVM is illustrated in Fig. 2. The system subscribes to IMU and LiDAR data simultaneously. All the IMU data received within the period of a LiDAR scan will be cached and propagated to the end of the scan to obtain a prediction of the state prior. Along with the estimated state and cached IMU data, LiDAR data will be fed into the LiDAR process module, where the points will be deskewed and projected to the end time of the LiDAR scan. Then each point is fitted into a Gaussian distribution with its neighbor points; thereafter, the points are projected to the global coordinate to find the corresponding voxels in the global map. Only the correspondences with similarity exceeding a certain threshold will be admitted in this procedure. The observation function will be formulated with all the accepted correspondences and fed into the IESKF state estimator for the estimation of the current state. The Incremental Map module consists of nothing but a simple, consistently updated global voxel map. Given the estimate of the last scan: $\bar{\mathbf{x}}_{k-1}$ and $\bar{\mathbf{P}}_{k-1}$, we will detail the system workflow in turn in the following subsections.
@@ IV-A IMU Process
Since the sampling period of IMU is much shorter than that of the LiDAR scan, the state prediction will be performed recursively until the end time of the current scan $t_{k}$. Recalling the state defined in (1), we have the following discrete model based on continuous kinematics equations [25]:
@@
where $\mathbf{u}_{i}$, $\mathbf{w}_{i}$, and $\mathbf{f}$ are the system input (the IMU data), system noise, and kinematics function respectively. Please refer to Appendix. -A for the detailed formulations. Firstly, the mean of the nominal state $\hat{\mathbf{x}}_{i}$ will be predicted without considering the noise $\mathbf{w}_{i}$:
@@
$\delta\mathbf{x}_{i+1}\approx\mathbf{F}_{\delta\mathbf{x}}\delta\mathbf{x}_{i}+\mathbf{F}_{\mathbf{w}}\mathbf{w}_{i};\delta\mathbf{x}_{0}=\delta\bar{\mathbf{x}}_{k-1}=\mathbf{0}.\$ (5)
Readers may refer to [10] for the derivation and detailed formulation of $\mathbf{F}_{\delta\mathbf{x}}$ and $\mathbf{F}_{\mathbf{w}}$. In implementation, the system noise $\mathbf{w}_{i}$ is considered as Gaussian white noise: $\mathcal{N}(\mathbf{0},\mathbf{W})$. The error state $\delta\mathbf{x}_{i}$ is independent random vector with multivariate normal distribution: $\delta\mathbf{x}_{i}\sim\mathcal{N}\left(\mathbf{0},\mathbf{P}_{i}\right)$. As a consequence, the predicted covariance matrix $\hat{\mathbf{P}}_{i+1}$ of $\delta\mathbf{x}_{i+1}$ can be calculated as the linear transformation:
@@ IV-B LiDAR Process
All the LiDAR points ${}^{L_{j}}\mathbf{p}_{j}$ with timestamps $\gamma_{j}\in\left[t_{k-1},t_{k}\right]$ are cached to form a point cloud $\mathcal{P}_{k}$. ${}^{L_{j}}\mathbf{p}_{j}$ represents the point with coordinate $\mathbf{p}$ under the LiDAR frame $L_{j}$ sampling at time $\gamma_{j}$. For most commercial LiDAR, the points are sampled under different timestamps, which will introduce motion distortion to the cloud $\mathcal{P}_{k}$. We seek to compensate for the distortion by projecting all the points in $\mathcal{P}_{k}$ to $L_{k}$, the LiDAR frame at $t_{k}$. To this end, We adopt the backward propagation proposed in [9] to obtain the relative motion between IMU frames sampled at $\gamma_{j}$ and $t_{k}$: ${}^{I_{k}}\mathbf{T}_{I_{j}}$. The point ${}^{L_{j}}\mathbf{p}_{j}$ can thus be projected to $L_{k}$ as:
@@
Moreover, a representative descriptor for the point is required for robust correspondence matching. Thus, we represent the deskewd points as a collection of Gaussian distributions. For each point ${}^{L_{k}}\mathbf{p}_{j}$, we collect its nearest $y$ neighbors as $\{\mathbf{a}_{i}\}$ with $i=0,1,...,y$ and assume they are sampled from a Gaussian distribution $\mathcal{N}({}^{L_{k}}\boldsymbol{\mu}_{j},{}^{L_{k}}\mathbf{C}_{j})$. The maximum likelihood estimate of the distribution is:
${}^{L_{k}}\boldsymbol{\mu}_{j}$ $=\frac{1}{y}\sum_{i=0}^{y}\mathbf{a}_{i},$ (8)
$\mathbf{A}$ $=\begin{bmatrix}\mathbf{a}_{0}-{}^{L_{k}}\boldsymbol{\mu}_{j}&...&\mathbf{a}_{y}-{}^{L_{k}}\boldsymbol{\mu}_{j}\end{bmatrix},$
${}^{L_{k}}\mathbf{C}_{j}$ $=\frac{1}{y-1}\mathbf{A}\mathbf{A}^{T}.$
The Gaussian distribution assumption can be considered a good estimate of the local surface shape at a sufficiently small scale [11]. Eventually, the LiDAR process module outputs a collection of the fitted Gaussian distributions for all the points: $\{{}^{L_{k}}\boldsymbol{\mu}_{j},{}^{L_{k}}\mathbf{C}_{j}\}$.

### v3

@@ Abstract
This letter presents a probabilistic voxel-based LiDAR Inertial Odometry framework for accurate and robust pose estimation. The framework addresses the correspondence mismatching issue by representing the LiDAR points as a set of Gaussian distributions and evaluating the divergence in variance for outlier rejection. Based on the fitted distributions, a new residual metric is proposed for the filter-based Lidar inertial odometry by incorporating both the distance and variance disparities, further enriching the comprehensiveness and accuracy of the residual metric. With the strategic design of the residual, we propose a simple yet effective voxel-solely mapping scheme, which only requires the maintenance of one centroid and one covariance matrix for each voxel. Experiments on different datasets demonstrate the robustness and high accuracy of our framework for various data inputs and environments. To the benefit of the robotics society, we open-source the code at https://github.com/Ji1Xingyu/lio_gvm.
@@
LiDAR inertial odometry, SLAM, probabilistic feature association, voxel map
@@ I Introduction
The vital requirement for an autonomous robot is to estimate its ego-motion and build a map of the environment without prior knowledge, which is widely known as Simultaneous Localization and Mapping (SLAM). In particular, the LiDAR and IMU sensors are commonly employed together [1, 2, 3, 4, 5, 6] in SLAM problem to form a LiDAR Inertial Odometry (LIO) because of their complementary ability: LiDAR contributes accurate range measurements for the environment, while IMU provides precise and surrounding-insensitive short-term motion information [7]. Because of the temporal and computational efficiency, filter-based LIO [3, 4, 5, 6] has been investigated in the recent few years. The performance of a LiDAR odometry is predominantly influenced by two components: point cloud registration and mapping scheme [8]. In this letter, we aim to enhance the performance of a filter-based LIO system by addressing improvements to the shortcomings of the existing works in these two areas.
Figure 1: The top image illustrates the global map built by LIO-GVM aligned with the Google map. The bottom image shows the runtime voxel map of LIO-GVM inside the rectangle. For better visualization, we model the distributions within each voxel as surfels. The blue, red, green, and white surfels denote the ground, wall, pole, and other types of voxels respectively.
Point cloud registration refers to aligning the source LiDAR scan with a target map together, namely, to estimate the optimal transformation between their local coordinates. Early works extract the geometric feature points from the input scan and match them with the nearest point in the target map [9, 10, 11]. A limitation of these methods is their deterministic target map representation, which cannot account for measurement uncertainties, resulting in inaccuracies in mapping and pose estimation. To address this issue, subsequent works model the target map in a probabilistic way, such as planes [12] or surfels [13, 14, 15]. The matched correspondences with large uncertainty will be depressed in the pose estimation phase. Nevertheless, they still employ the nearest correspondence matching. The key limitation of this matching scheme is its inability to reject false positive correspondences, which occurs when the source point is matched with the wrong target feature but identified as a valid pair. This issue can negatively affect the pose estimation accuracy.
The mapping scheme acts as the interdependent component with point cloud registration, which maintains the target map for efficient environment representation. A common method is to maintain the target map with a k-d tree [9, 10, 11] as it can provide strict K nearest neighbor (k-NN) search for the correspondence matching. The time complexity for the deletion, insertion, and query operations of the k-d tree is $\mathcal{O}(logN)$, which can impact real-time performance when dealing with a significantly large target map. To address this problem, some systems opt to govern the target map with hash-based voxels [5, 12, 16], with $\mathcal{O}(1)$ the time complexity for element operation. However, due to the requirement of the front end, these voxel-based maps need to embed a supplemental data structure (k-d tree, Hilbert Curve, etc.) within voxels, which cannot reach the optimal operation speed.
To cope with the problem of mismatch in point cloud registration, we represent the points in both the source and target clouds as Gaussian distributions [17, 18, 19] and evaluate the matched correspondences by the disparity in variance. Moreover, based on the distribution assumption, we propose a new residual metric for filter-based LIO. The formula accesses both the distance and variance disparities, further enriching the comprehensiveness and accuracy of the residual metric. For the mapping side, instead of storing all the points, we propose an incremental voxel map without any additional structures, which only maintains a centroid and covariance matrix for each voxel and achieves $\mathcal{O}(1)$ time complexity.
@@
In conclusion, we implement the proposed point cloud registration and mapping scheme to a filter-based LIO system denoted as LIO-GVM. The contributions can be summarized as follows:
  • We propose a new correspondence matching method and residual metric for filter-based LIO. This matching scheme can reject false positive correspondences. Additionally, the metric measures not only distance but also variance disparity, enhancing the comprehensiveness and accuracy of the residual formulation.
@@
  • We propose a simple yet efficient mapping scheme. The map only necessitates the maintenance of the centroid and covariance matrix for each voxel, which can achieve $\mathcal{O}(1)$ time complexity for element operation.
@@
  • We implement the two key techniques into a filter-based LIO system and evaluate its robustness and efficiency through various experiments (see Fig. 1). To the benefit of the robotics community, our implementation code is open-sourced at https://github.com/Ji1Xingyu/lio_gvm.
II Related Work
For point cloud registration, LOAM [9] first proposes to extract planar and edge features from the source and target clouds for fast registration and achieves accurate results. LeGO-LOAM [10] performs Euclidean cluster first to extract the ground plane and restrict the correspondence match in the same cluster. Instead of extracting the geometric features from the source scan, some systems directly register the raw source cloud into the target map [4, 3]. SuMa [13] and SLICT [14] render points in the target map into surfels and consider the point-to-surfel distance as residual. Similarly, VoxelMap [12] proposes to maintain a probabilistic plane-based target map and register the point in the source scan to the map. These point-to-surfel (or plane) residuals consider the noise in the target map and thus can depress the influence of matched pairs with high uncertainty in the pose estimation phase. Nevertheless, they cannot handle the false positive correspondence problem. On the contrary, our proposed matching scheme can evaluate the similarity of the matched distributions and reject the false positive pairs.
The proposed residual metric is a weighted distribution-to-distribution (D2D) distance and is similar to the D2D residual in GICP [18] and D2D-NDT [20]. The key distinction lies in the proposed similarity-based weighting scheme, which can adjust the weight of the residual for matched correspondences based on their divergence in covariance. Subsequently, the proposed residual metric can provide a more comprehensive and accurate pose estimation. Similarly, LiTAMIN2 [21] introduces symmetric KL-Divergence to the residual metric for divergence evaluation. However, their proposed formulation cannot be directly deployed in filter-based pose estimation.
For mapping schemes, a common method is to maintain the target map with a k-d tree [9, 11]. The conventional k-d tree needs to rebuild the entire tree for the map update, which is inefficient and time-consuming. FAST-LIO2 [4] uses the ikd-Tree [22], a dynamic k-d tree that can incrementally update the map. VoxelMap, Faster-LIO [5], and SuperOdom [16] govern the target map with hash-based voxels. However, due to the requirement of the front end, these voxel-based mapping schemes need to embed a supplemental data structure within voxels to form a hierarchical structure for the maintenance of the raw points. Instead, the proposed mapping scheme only requires the maintenance of the centroid and covariance matrix for each voxel, resulting in a time complexity of $\mathcal{O}(1)$ for element operation.
@@ III Preliminary
${}^{L_{k}}\mathbf{p}_{j}$ The $j_{th}$ point coordinate with reference to $L$ at time $t_{k}$
${}^{G}\mathbf{v}_{i}$ The coordinate of the $i_{th}$ voxel with reference to $G$
@@ III-A Notation
The important notations used in this letter are listed in Tab. I. Besides, we assume a static calibrated extrinsic matrix ${}^{L}\mathbf{T}_{I}=({}^{L}\mathbf{R}_{I},{}^{L}\mathbf{p}_{I})$ from IMU frame to LiDAR frame.
@@ III-B System Description
where ${}^{G}\mathbf{v}_{I}$ and ${}^{G}\mathbf{g}$ are the velocity of IMU and gravity vector with reference to the global frame, $\mathbf{b}_{\boldsymbol{\omega}}$ and $\mathbf{b}_{\mathbf{a}}$ are the bias vectors of gyroscope and accelerometer of the IMU. The true state $\mathbf{x}$ of the system can be divided into two states:
@@
$\delta\mathbf{x}_{k}$ $=\mathbf{x}_{k}\boxminus\hat{\mathbf{x}}_{k}=\begin{bmatrix}{}^{G}\delta\mathbf{r}_{I_{k}}^{T}&{}^{G}\delta\mathbf{p}_{I_{k}}^{T}&{}^{G}\delta\mathbf{v}^{T}_{I_{k}}&\delta\mathbf{b}^{T}_{\boldsymbol{\omega}_{k}}&\delta\mathbf{b}^{T}_{\mathbf{a}_{k}}&{}^{G}\delta\mathbf{g}_{k}^{T}\end{bmatrix}^{T},$
@@ IV Methodology
Given the estimate of the last scan: $\bar{\mathbf{x}}_{k-1}$ and $\bar{\mathbf{P}}_{k-1}$, we will detail the system workflow in the following subsections.
@@ IV-A IMU Process
Recalling the state defined in (1), we have the following discrete model based on continuous kinematics equations [23]:
@@
where $\mathbf{u}_{i}$, $\mathbf{w}_{i}$, and $\mathbf{f}$ are the system input, system noise, and kinematics function respectively. Please refer to Appendix -A for the detailed formulations. Firstly, the mean of the nominal state $\hat{\mathbf{x}}_{i}$ will be predicted without considering the noise $\mathbf{w}_{i}$:
@@
$\delta\mathbf{x}_{i+1}\approx\mathbf{F}_{\delta\mathbf{x}}\delta\mathbf{x}_{i}+\mathbf{F}_{\mathbf{w}}\mathbf{w}_{i};\delta\mathbf{x}_{0}=\delta\bar{\mathbf{x}}_{k-1}=\mathbf{0},\$ (5)
where $\mathbf{F}_{\delta\mathbf{x}}$ and $\mathbf{F}_{\mathbf{w}}$ are the Jacobian matrices of $\mathbf{f}()$ w.r.t. $\delta\mathbf{x}_{i}$ and $\mathbf{w}_{i}$ [4]. In the implementation, the system noise $\mathbf{w}_{i}$ is considered as Gaussian white noise following the normal distribution $\mathcal{N}(\mathbf{0},\mathbf{W})$, where $\mathbf{W}$ is the covariance matrix. The error state $\delta\mathbf{x}_{i}$ is an independent random vector with multivariate normal distribution: $\delta\mathbf{x}_{i}\sim\mathcal{N}\left(\mathbf{0},\mathbf{P}_{i}\right)$. As a consequence, the predicted covariance matrix $\hat{\mathbf{P}}_{i+1}$ of $\delta\mathbf{x}_{i+1}$ can be calculated as the linear transformation:
@@ IV-B LiDAR Process
All the LiDAR points ${}^{L_{j}}\mathbf{p}_{j}$ with timestamps $\gamma_{j}\in\left[t_{k-1},t_{k}\right]$ are cached to form a point cloud $\mathcal{P}_{k}$. For most commercial LiDAR, the points are sampled under different timestamps, which will introduce motion distortion to the cloud $\mathcal{P}_{k}$. We seek to compensate for the distortion by projecting all the points in $\mathcal{P}_{k}$ to $L_{k}$, the LiDAR frame at $t_{k}$. To this end, we adopt the backward propagation proposed in [3] to obtain the relative motion between IMU frames sampled at $\gamma_{j}$ and $t_{k}$: ${}^{I_{k}}\mathbf{T}_{I_{j}}$. The point ${}^{L_{j}}\mathbf{p}_{j}$ can thus be projected to $L_{k}$ as:
@@
Moreover, a representative descriptor for the point is required for robust correspondence matching. Thus, we represent the deskewd points as a collection of Gaussian distributions. For each point ${}^{L_{k}}\mathbf{p}_{j}$, we collect its nearest $y$ neighbors and fit them to a Gaussian distribution $\mathcal{N}({}^{L_{k}}\boldsymbol{\mu}_{j},{}^{L_{k}}\mathbf{C}_{j})$ [17]. Eventually, the LiDAR process module outputs a collection of the fitted Gaussian distributions for all the points: $\{{}^{L_{k}}\boldsymbol{\mu}_{j},{}^{L_{k}}\mathbf{C}_{j}\}$.